# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

### v2

@@ IV-D1 Static Weighted Consolidation
IV-D2 Adaptive Weighted Consolidation
@@
By applying Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
@@
Skill consolidation is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization. Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$. Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
Algorithm 2 TaSL
@@
V TaSL with Memory Replay: TaSL-M
To adapt TaSL for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, TaSL-M, designed to further improve model performance. The implementation process is shown in Figure 3.
Figure 3: Overview of TaSL-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of TaSL.
Following the training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$. Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$. This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge.
@@ VI-A1 Datasets
Further details about the tasks and their order are provided in the Appendix.
@@
TaSL (ours) 26.41 11.78 -18.55 70.71 11.80 -3.27
TasLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -2.04
Replay 35.37 2.35 -15.79 75.28 3.28 -1.88
@@
TaSL-M (ours) 52.12 11.13 -1.31 84.26 12.32 -1.08
TasLoRA-M (ours) 53.01 11.21 -0.81 84.33 11.71 -0.98
@@ VI-A3 Baselines
We evaluate TaSL against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders. IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks. Replay: replays real samples from old tasks when learning new tasks to avoid forgetting. EWC [23]: finetune the model with a regularization loss designed to minimize updates to parameters critical to previously learned tasks. L2P [42]: uses the input to dynamically select and update prompts from a fixed prompt pool. LFPT5 [43]: continuously trains a soft prompt for each task with generative replay. Prog-Prompt [17]: sequentially concatenates previous learned prompts to the current one during the training and testing time. O-LoRA [18]: learns tasks in distinct LoRA subspaces, ensuring orthogonality, and aggregates all LoRA weights for testing. SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
Table II compares our TaSL with these baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities. To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for skill localization (Eq. 8), and employs static weighted consolidation (Eq. 15). TasLoRA: as the most recent extension of TaSL, this variant introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric for skill localization (Eq. 9), and features adaptive weighted consolidation (Eq. 17). TasL-M and TasLoRA-M: these enhancements of TaSL and TasLoRA integrate memory replay to further boost performance.
TABLE II: The comparison between TaSL and other CL methods. Specifically, RF indicates whether the method is rehearsal-free. PE indicates whether the method is parameter efficient. UT indicates whether the method can be applied to solve unseen tasks. KT indicates whether the method enables knowledge transfer.
@@ VI-A4 Implementation Details
We utilize two distinct language model architectures previously adopted in the field of CL for NLP: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50]. In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (10) and Eq. (11) are set to 0.85. We set $\beta$ in Eq. (13) to 0.7 and $\tau$ in Eq. (16) to 0.15. Following [51], we allocate 2% of the original training set volume for replay samples in all replay-based baseline methods. For different backbones, we utilized the following hyperparameters:
@@ VI-B Main Results
Firstly, TaSL and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer. Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
Furthermore, our enhanced version, TasLoRA, significantly outperforms TaSL. Specifically, TasLoRA achieves average gains of 1.3% in AP and 1.4% in BWT, underscoring the effectiveness of enhancements to each framework component. Notably, even without memory replay, TasLoRA nearly matches the performance of SAPT with memory replay on the Long Sequence Benchmark, particularly in BWT, with only a minor difference of 0.8% (-2.04% vs. -1.25%).
For memory-based methods, both TaSL-M and TasLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework. For subsequent experimental analyses and comparisons, we will focus on TasLoRA, the most effective memory-free version, for further evaluations.
Secondly, TasLoRA consistently demonstrates superior performance across various backbones. To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones. As depicted in Figure 4, performance improvements are evident with increasing model size. For instance, in T5-XL, TasLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively. These results further validate the robust generalization ability of our method.
@@
Figure 4: Performance of TasLoRA across various backbones.
Thirdly, our skill consolidation technique effectively counters catastrophic forgetting. To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark. Figure 5 illustrates that TasLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task. In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
@@ VI-C Performance on Unseen Tasks
Building on prior research [15], we selected three tasks from each category to evaluate our method’s cross-task generalization capabilities, a critical measure for CL algorithms. The results, indicated by the average Rouge-L scores in Table III, highlight the effectiveness of our approach. T5-ZS refers to the zero-shot task adaptation methods employed. TasLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
@@ VI-D Visualization of Skill Units
  • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters. This observation substantiates the feasibility of the design motivations behind the TaSL framework.
@@
TABLE IV: Ablation Study for Importance-Aware Skill Localization.
@@ VI-E Ablation Study
In this section, we assess the impact of importance-aware skill localization, fine-grained skill consolidation, and memory size on model performance. Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
VI-E1 Effect of the Proposed Importance Metric in Skill Localization
We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq. (6) to focus solely on sensitivity, as outlined in Eq. (9); (ii) To verify the validity of our proposed approximate second-order gradient importance metric in Eq. (6), we implement a first-order Taylor expansion as described in Eq. (8). The comparative results, presented in Table IV, show that using moving averages for importance scoring substantially enhances model performance. Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics. These findings underscore the significant role of accurate skill localization in elevating model effectiveness.
VI-E2 Effect of Adaptive Skill Consolidation
We evaluate our fine-grained, skill unit-level adaptive weighted consolidation against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq. (14), and (ii) Exponential Moving Average (EMA) [52], which computes a running average of parameters at each fine-tuning iteration.
@@ VI-E2 Effect of Adaptive Skill Consolidation
TABLE V: Ablation Study for Fine-Grained Skill Consolidation.
@@ VI-E3 The Effect of Memory Size
We examine how varying the memory size affects the performance of Replay, TaSL-M, and TasLoRA-M. By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI. As expected, increasing the memory size generally enhances the performance of all methods, with Replay showing the most significant improvement. This is predictable as Replay heavily depends on memory to mitigate catastrophic forgetting. When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
Conversely, TaSL-M and TasLoRA-M utilize skill consolidation strategies to effectively maintain parameters that store historical knowledge. Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
@@ VII Conclusion
In this paper, we introduced the novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. TaSL employs importance-aware skill localization and fine-grained skill consolidation to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer. The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance. Extensive testing shows that TaSL excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches. Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
Despite significant improvements in efficiency and performance in large language model continual learning, certain limitations persist. For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of skill consolidation. Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance. Additionally, merging the skill localization and consolidation phases could allow for ongoing parameter consolidation based on their importance during model training. This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.

### v3

@@ IV-D1 Static Weighted Fusion
IV-D2 Adaptive Weighted Fusion
@@
By applying Eq. (15) or Eq. (17) in our knowledge fusion phrase, it effectively addresses the challenges in CL:
@@
Knowledge fusion is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization. Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$. Detailed implementation of entire KlF algorithm is provided in Algorithm 2.
Algorithm 2 KlF
@@
V KlF with Memory Replay: KlF-M
To adapt KlF for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, KlF-M, designed to further improve model performance. The implementation process is shown in Figure 3.
Figure 3: Overview of KlF-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of KlF.
Following the training on task $\mathcal{T}_{k}$ using the knowledge localization and fusion, we obtain the model $\hat{f}_{k}$. Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$. This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge.
@@ VI-A1 Datasets
KlF (ours) 26.41 11.78 -18.55 70.71 11.80 -3.67
KlFLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -3.04
Replay 35.37 2.35 -15.79 70.98 3.28 -15.42
@@
KlF-M (ours) 52.12 11.13 -1.31 84.21 12.32 -1.08
KlFLoRA-M (ours) 53.01 11.21 -0.81 84.36 11.71 -0.98
@@ VI-A3 Baselines
We evaluate KlF against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders. IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks. Replay: replays real samples from old tasks when learning new tasks to avoid forgetting. EWC [23]: finetune the model with a regularization loss designed to minimize updates to parameters critical to previously learned tasks. L2P [42]: uses the input to dynamically select and update prompts from a fixed prompt pool. LFPT5 [43]: continuously trains a soft prompt for each task with generative replay. Prog-Prompt [17]: sequentially concatenates previous learned prompts to the current one during the training and testing time. O-LoRA [18]: learns tasks in distinct LoRA subspaces, ensuring orthogonality, and aggregates all LoRA weights for testing. SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
Table II compares our KlF with other baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities. To clarify the specific localization and fusion techniques used in KlF and its variants: KlF: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for knowledge localization (Eq. 8), and employs static weighted fusion (Eq. 15). KlFLoRA: an extension of KlF that introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric (Eq. 9), and features adaptive weighted fusion (Eq. 17). KlF-M and KlFLoRA-M: these enhancements of KlF and KlFLoRA integrate memory replay to further boost performance.
TABLE II: The comparison between KlF and other CL methods. Specifically, RF indicates whether the method is rehearsal-free. PE indicates whether the method is parameter efficient. UT indicates whether the method can be applied to solve unseen tasks. KT indicates whether the method enables knowledge transfer.
@@ VI-A4 Implementation Details
We utilize two distinct language model architectures: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50]. In KlF, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (10) and Eq. (11) are set to 0.85. We set $\beta$ in Eq. (13) to 0.7 and $\tau$ in Eq. (16) to 0.15. Following [51], we allocate 2% of the original training set volume for replay samples in all replay-based baseline methods. Hyperparameters are as follows:
@@ VI-B Main Results
Firstly, KlF and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer. Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
Furthermore, our enhanced version, KlFLoRA, significantly outperforms KlF. Specifically, KlF-M achieves average gains of 1.3% in AP and 0.3% in BWT, underscoring the effectiveness of enhancements to each framework component. Notably, even without memory replay, KlF-M surpasses the performance of vanilla repla on the Long Sequence Benchmark, particularly in BWT.
For the memory-based methods, to ensure a fair comparison, we adopted the same data synthesis-based generative replay approach used in SAPT, ensuring that the amount of data used in memory replay is consistent across methods. The results show that both KlF-M and KlFLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework. For subsequent experimental analyses and comparisons, we will focus on KlFLoRA, the most effective memory-free version, for further evaluations.
Secondly, KlFLoRA consistently demonstrates superior performance across various backbones. To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones. As depicted in Figure 4, performance improvements are evident with increasing model size. For instance, in T5-XL, KlFLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively. These results further validate the robust generalization ability of our method.
@@
Figure 4: Performance of KlFLoRA across various backbones.
Thirdly, our knowledge fusion technique effectively counters catastrophic forgetting. To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark. Figure 5 illustrates that KlFLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task. In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
@@ VI-C Performance on Unseen Tasks
Building on prior research [15], we selected three tasks from each category to evaluate our method’s cross-task generalization capabilities, a critical measure for CL algorithms. The results, indicated by the average Rouge-L scores in Table III, highlight the effectiveness of our approach. T5-ZS refers to the zero-shot task adaptation methods employed. KlFLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
@@ VI-D Visualization of Skill Units
  • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters. This observation substantiates the feasibility of the design motivations behind the KlF framework.
@@
TABLE IV: Ablation Study for Importance-Aware Knowledge Localization.
@@ VI-E Ablation Study
In this section, we assess the impact of importance-aware knowledge localization, fine-grained knowledge fusion, and memory size on model performance. Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
VI-E1 Effect of the Proposed Importance Metric in Knowledge Localization
We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq. (6) to focus solely on sensitivity, as outlined in Eq. (9); (ii) To verify the validity of our proposed approximate second-order gradient importance metric in Eq. (6), we implement a first-order Taylor expansion as described in Eq. (8). The comparative results, presented in Table IV, show that using moving averages for importance scoring substantially enhances model performance. Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics. These findings underscore the significant role of accurate knowledge localization.
VI-E2 Effect of Adaptive Knowledge Fusion
We evaluate our fine-grained, skill unit-level adaptive weighted fusion against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq. (14), and (ii) Exponential Moving Average (EMA) [52], which computes a running average of parameters at each fine-tuning iteration.
@@ VI-E2 Effect of Adaptive Knowledge Fusion
TABLE V: Ablation Study for Fine-Grained Knowledge Fusion.
@@ VI-E3 The Effect of Memory Size
We examine how varying the memory size affects the performance of Replay, KlF-M, and KlFLoRA-M. By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI. As expected, increasing the memory size generally enhances the performance of all methods, with Replay showing the most significant improvement. This is predictable as Replay heavily depends on memory to mitigate catastrophic forgetting. When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
Conversely, KlF-M and KlFLoRA-M utilize knowledge fusion strategies to effectively maintain parameters that store historical knowledge. Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
@@ VII Conclusion
In this paper, we introduced the novel Knowledge Localization and Fusion (KlF) framework for language model continual learning. KlFLoRA employs importance-aware knowledge localization and fine-grained knowledge fusion to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer. The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance. Extensive testing shows that KlFLoRA excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches. Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
Despite significant improvements in efficiency and performance in large language model continual learning, certain limitations persist. For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of knowledge fusion. Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance. Additionally, merging the knowledge localization and fusion phases could allow for ongoing parameter consolidation based on their importance during model training. This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.